# Understanding the Challenges and Promises of Developing Generative AI Apps: An Empirical Study

### v2

@@ 2.2. User Review Analysis Techniques
In this section, we summarize prior work related to 1) filtering user reviews and 2) topic extraction and assignment to user reviews.
@@
Filtering user reviews. Automated filtering has shown strong potential in enhancing user review analysis by effectively removing non-informative content. Chen et al. (Chen et al., 2014) introduced a framework called AR-miner that used Latent Dirichlet Allocation (LDA) to first filter informative reviews with a hit-rate of 70%, and second to group the informative reviews using topic modelling. Ghosh et al. (Ghosh et al., 2024) evaluated fine-tuned LLMs for filtering non-informative user reviews on the Google Play Store and found them highly effective, achieving accuracies of 92% for both BERT and DistilBERT, and 91% for Gemma. Al Wahshat et al. (Al Wahshat et al., 2023) highlighted the potential of using GPT-4 in filtering manipulated reviews due to its powerful language processing capabilities.
@@
Topic extraction and assignment to user reviews. Before the advancement of LLMs, researchers used various methods to analyze user reviews. Genc Nayebi and Abran (Genc-Nayebi and Abran, 2017) conducted a literature review in 2016 that explored approaches for mining opinions from user reviews. They identified three studies that relied on manual analysis, while others used automated techniques, such as Classification and Regression Trees (CART), word frequency statistics, correlation analysis, LDA, and Hexbin plots. Researchers and developers still use non-LLM approaches, even in recent work. For example, Arambepola et al. (Arambepola et al., 2024) used the Valence Aware Dictionary and sEntiment Reasoner (VADER) for sentiment analysis to categorize reviews into positive, neutral, and negative categories. They then used an LDA model for topic modelling of user reviews of educational apps. Their positive topics achieved a coherence score of 0.56, and their negative topics achieved a score of 0.49. Assi et al. (Assi et al., 2021a) proposed the Global-Local sensitive Feature Extractor (GLFE) to identify high-level features in user reviews, achieving 79–82% precision and 74–77% recall on annotated review data.
@@
LLMs have demonstrated promising performance in tasks related to topic extraction and assignment. Arambepola et al. (Arambepola et al., 2025) conducted a recent systematic review of mobile app review analysis techniques and found that integrating LLMs into app user review analysis provides a more holistic and context-rich approach to evaluating reviews and understanding users. Pham et al. (Pham et al., 2024) introduced TopicGPT, a framework to perform topic extraction by prompting Open AI’s GPT (OpenAI, 2023) to generate topics and assign them to documents. Their framework outperformed traditional topic models, such as LDA, SeededLDA, and BERTopic. Prakash et al. (Prakash et al., 2023) introduced PromptMTopics, a prompt-based ChatGPT model set to extract topics from textual meme captions that outperformed baseline models. Assi et al. (Assi et al., 2025a) introduced LLM-Cure, an approach that extracts and assigns features from user reviews to make suggestions for feature improvement. Their approach outperformed baseline models, achieving an F1-score of 85%. Sorathiya and Ginde (Sorathiya and Ginde, 2024) employed a hybrid approach that combines Natural Language Inference (NLI) and LLMs to identify privacy concerns in mental health app reviews, aiming to uncover ethically relevant user feedback that traditional keyword-based methods often overlook. Their approach achieved an F1-score of 81%.
@@
Instead of extracting and assigning topics to user reviews, prior studies focused on a single task (i.e., either extracting topics from user reviews or assigning predefined topics to user reviews). For example, Wei et al. (Wei et al., 2023) focused on review assignment using predefined labels rather than extracting topics. They assigned reviews using fixed categories such as feature requests, problem reports, and irrelevant feedback. While useful, this approach limits the discovery of emergent or evolving user concerns. Dos Santos et al. (Dos Santos et al., 2023) extracted topics from user reviews without performing the assignment task. Their study examined the impact of app updates on accessibility by utilizing GPT-4 to identify key issues, barriers, improvements, update consequences, and user sentiments. While this approach is effective for uncovering topical insights, it does not account for the frequency or distribution of each topic, limiting its usefulness for trend or prevalence analysis. Ren et al. (Ren et al., 2024) used LDA for topic extraction and pretrained LLMs for requirement extraction. Rezaei Nasab et al. (Rezaei Nasab et al., 2025) conducted a study of AI-based mobile app reviews to identify fairness-related concerns. They constructed a labelled dataset of fairness vs. non-fairness reviews, developed machine learning and LLM-based classifiers to detect fairness mentions, applied K-Means clustering, and performed manual coding to identify six types of fairness concerns.
Our work builds on this line of research by utilizing LLMs to filter out non-informative reviews and identify topics in user reviews of Gen-AI apps. Based on the analysis of user reviews, we identify the challenges and potential opportunities to adopt Gen-AI functionalities in mobile apps.
2.3. Prompt Design and Evaluation
Prompt engineering is the process of designing and refining prompts to effectively guide LLMs in performing specific tasks (Koyuncu, 2025). Prior work discusses prompt design in three directions: 1) the process of prompt design, 2) the definition of the required tasks, and 3) identifying the number of included examples.
The process of prompt design. Marvin et al. (Marvin et al., 2024) emphasized the importance of prompt refinement and outlined several prompt engineering techniques, including dynamic prompting, which refers to iteratively adjusting the prompt based on the generated output. He et al. (He et al., 2024) investigated the impact of prompt format across various tasks and found that prompt structure significantly affects performance, with different formats performing best in different task types. In our work, we adopt a prompt format inspired by prior work (Assi et al., 2025a) and refine it to suit our study’s goals, employing dynamic prompting to enhance accuracy and relevance further.
The definition of the required tasks. To ensure accurate, ethical, and reliable AI outputs, effective prompt engineering requires understanding common pitfalls, such as ambiguity, bias, and lack of context (Giray, 2023). Prompts can be explicit, implicit, or creative, varying in the amount of guidance they offer to the AI. Explicit prompts provide clear instructions, while implicit and creative prompts allow more flexibility and imagination but may produce less predictable results (Data Science Horizons, 2023). Ho et al. (Ho et al., 2024) adopted a prompt classification framework consisting of the three types. The authors analyzed how different prompts affect AI-generated responses. Explicit prompts generate specific and fact based answers, implicit prompts produce broader and more flexible responses, while creative prompts encourage innovation and originality. In our work, we adopt an explicit prompting strategy to improve structure and reproducibility.
Identifying the number of included examples. Deshmukh et al. (Deshmukh et al., 2025) conducted a literature review of prompt engineering and discussed 16 LLM prompt patterns to enhance output accuracy. Their analysis highlighted the importance of clarity, specificity, defined output formats, and contextual examples to improve model accuracy. However, it did not address techniques for optimizing example selection (i.e., determining the optimal number of examples to include in the prompt). Complimenting this, Wang et al. (Wang et al., 2023) demonstrated that using similar examples and increasing the number to four enhances performance, while adding more than four examples yields only marginal gains. Their work focused on classification, multiple-choice, and generation tasks, rather than user review analysis tasks. Prakash et al. (Prakash et al., 2023) evaluated prompts with 2, 4, 6, and 8 examples, finding comparable topic coherence and diversity across configurations.
In this paper, we further extend this line of research by investigating how the number of in-context examples and the exclusion of noisy data affect topic extraction and assignment tasks for user review. Our work demonstrates that increasing the number of examples for topic extraction from 0-shot to 5-shot significantly improves the model’s accuracy. Our work also suggests that filtering out non-informative reviews enhances the accuracy.
@@ 3. Methodology
To study user feedback on Gen-AI apps, we design a four-phase methodology we call SARA, which stands for Selection, Acquisition, Refinement, and Analysis. SARA encompasses 1) the selection of relevant Gen-AI apps, 2) the acquisition of user reviews, 3) a refinement data cleaning pipeline to ensure data quality for LLM analysis, and finally, 4) the LLM-based analysis of user reviews through topic extraction and assignment. Our proposed methodology enables us to address our RQs concerning the impact of data quality and few-shot examples on LLM performance, the prevalent topics within Gen-AI user reviews, and the temporal evolution of user ratings toward these apps. We present an overview of our research approach in Figure 2.
Figure 2. Overview of our approach
@@ 3.1. Selection of Gen-AI apps
    Keyword generation. We compile a list of keywords derived from a Gen-AI taxonomy by Gozalo-Brizuela and Garrido-Merchán (Gozalo-Brizuela and Garrido-Merchán, 2023). The authors organized Gen-AI apps into 14 categories, each with a taxonomy of the relevant technology. We exclude four categories, i.e.,“biotech,” “brain,” “drug discovery,” and “other,” because an Android mobile app user does not typically use them. We create a list of ten categories relevant to mobile apps and repurpose the taxonomy as keywords, as shown in Table 1.
@@
    Manual screening and validation. The initial search yields 303 apps. To ensure the inclusion of only genuine Gen-AI apps, the first and second authors independently review the apps by looking at the title, description, and images of the apps. We assess the inter-rater reliability, using Cohen’s Kappa (Cohen, 1960) ($\kappa=0.61$), which indicates substantial agreement. In case of disagreement, the two authors discuss and resolve the discrepancies for the ambiguous cases. This rigorous process results in a final list of 194 confirmed Gen-AI apps.
@@ 3.2. Acquisition of User Reviews
Following app selection, we collect user reviews using the Google Play Store Scraper (Mingyu, nd). For each retrieved review, the crawler extracts a comprehensive set of metadata, including reviewId, userName, userImage, content (i.e., review text), score (i.e., user rating from 1 to 5 stars), thumbsUpCount, reviewCreatedVersion, at (i.e., timestamp of the review), replyContent, repliedAt, and appVersion. Our subsequent analysis primarily focuses on the content, score, and at fields. We organize the collected reviews based on the Google Play Store categories (app categories) of their respective apps. At the end of this step, we collect 2,778,643 reviews of 186 apps, as eight apps from our list have no textual reviews.
@@ 3.3. Multi-Stage User Review Refinement
To ensure the robustness and reliability of our LLM-based analysis, we implement a multi-stage data cleaning pipeline designed to enhance the quality of the user review data progressively. This multi-stage refinement process comprises four distinct stages. We refine data in stages because we are interested in experimenting with different levels of data cleaning in RQ1. We summarize the number of reviews retained after each cleaning stage in Table 2. Each stage builds upon the previous one as follows:
Table 2. Number of reviews by app category and stage.
@@
Stage 0: basic cleaning. Text preprocessing, as demonstrated in prior work (Ghosh et al., 2024; Prakash et al., 2023; Roumeliotis et al., 2024; Lee et al., 2024b), is a standard practice for eliminating noise and ensuring more effective analysis. It involves the removal of punctuation, special characters, extra white spaces, emojis, and irrelevant data such as usernames and URLs. In the initial stage, we exclude empty reviews and apps with no reviews, as we are interested in analyzing the textual content of the reviews. Then, we perform fundamental text preprocessing, including replacing commas with white spaces and removing special characters and emojis from the review content.
@@
We exclude five categories that contain fewer than 1,000 reviews, as small sample sizes tend to produce repetitive and low-quality topics. To validate this threshold, we experiment with categories of varying sizes and find that categories with fewer than 1,000 reviews yield topics that lack diversity and informativeness. For instance, the Personalization app category, which contains only 23 reviews, produces two nearly identical and non-informative topics labelled Positive Feedback and Praise. Additionally, some topics focus on non-informative content, such as Personal Comments or Signatures, which capture examples like “hello marlyn” and “Placau!!!!” that are not relevant to the app’s functionality or performance. Based on these observations, we adopt 1,000 reviews as the minimum threshold and exclude any app categories that fall below this cutoff.
@@
Stage 1: temporal filtering. This stage filters reviews written prior to 2022 to ensure Gen-AI relevancy. Some apps included in our initial screening are not originally designed as Gen-AI tools but have integrated Gen-AI functionalities at a later stage. For instance, in the case of Microsoft Bing Search [^4] https://play.google.com/store/apps/details?id=com.microsoft.bing, Bing Chat is a recently introduced feature that extends the capabilities of the search engine by incorporating Gen-AI models (Iorliam and Ingio, 2024). Consequently, it is essential to exclude user reviews published before the integration of such features, as they reflect experiences with the app before the adoption of Gen-AI. While foundational work on AI models began as early as 2012, notable advancements in Gen-AI only started to gain momentum after 2018 (Sengar et al., 2024). The year 2018 also marked a rise in academic interest and publications related to Gen-AI (Dwivedi and Elluri, 2024). However, the field witnessed a major breakthrough in 2022, particularly with the public release and widespread adoption of ChatGPT, which significantly accelerated the prominence and deployment of Gen-AI technologies across various platforms (Akhtar, 2024). To maximize the relevance and accuracy of the user feedback analyzed in our study, we adopt 2022 as the cutoff year for including reviews.
@@
Stage 2: short review exclusion. Building upon the cleaned data from Stage 1, this stage employs a simple heuristic to further reduce noise by excluding reviews containing three words or less. Short reviews of three words or less do not contain context for meaningful topics (Assi et al., 2021a; Arambepola et al., 2024). This step aims to eliminate overly brief and potentially non-informative feedback.
@@
Stage 3: LLM-assisted informative review filtration. Non-informative reviews bring noise to user reviews analysis and influence the performance of topic extraction (Chen et al., 2014). To improve the quality of extracted topics, we exclude non-informative reviews. We define non-informative reviews as short, generic statements that express sentiment without referencing specific reasons or features. While Chen et al. (Chen et al., 2014) focused on informativeness from a developer’s perspective, we expand this definition to consider the informational needs of multiple stakeholders, including users, platform owners, developers and researchers. We carry out the LLM-assisted process in the following steps.
@@
    Prompt design. To filter non-informative reviews, we utilize Open AI’s (OpenAI, 2023) GPT-4o-mini model with zero temperature for reproducibility. We use gpt-4o-mini instead of the full gpt-4o model because it offers significantly faster response times while maintaining reasonable output quality for our task, which does not require complex reasoning (OpenAI, 2024b).
    We design a filtering prompt that instructs the LLM to classify reviews as either informative or non-informative. To guide the model effectively, we incorporate representative examples within the prompt.
@@
    Iterative prompt refinement. We use small statistically significant samples of 96 reviews (95% confidence interval, 10% margin of error) from three diverse app categories (Photography, Productivity, and Art & Design). Then, the first and second authors manually label the reviews in the samples as informative or non-informative. We calculate the inter-rater reliability using Kappa (Cohen, 1960) ($\kappa=0.91$), which indicates almost perfect agreement.
    Next, we design and test different versions of the filtering prompt against the manually labelled data, evaluating the model’s filtering accuracy and iteratively refining the prompt. Revisions include modifying the wording, structure, number of examples, and types of examples. We repeat this iterative process until we achieve a peak filtering accuracy of 91% on held-out validation samples. Figure 3 shows the final version of the filtering prompt, which includes ten examples (five informative and five non-informative) and yields optimal performance.

### v3

@@ 2.2. User Review Analysis Techniques
In this section, we summarize prior work related to 1) filtering user reviews, 2) pre-LLM approaches to topic extraction, 3) LLM-based topic extraction and assignment methods, 4) single-task topic analysis, and 5) domain-specific applications of LLMs in user review analysis. These studies collectively form the foundation for our own framework SARA, which builds upon their methods but advances them by integrating filtering, topic extraction, and topic assignment into a unified, scalable framework for analyzing Gen-AI app reviews.
@@
Filtering user reviews. Automated filtering has shown strong potential in enhancing user review analysis by effectively removing non-informative content. Chen et al. (Chen_Lin_Hoi_Xiao_Zhang_2014) introduced a framework called AR-miner that used Latent Dirichlet Allocation (LDA) to first filter informative reviews with a hit-rate of 70%, and second to group the informative reviews using topic modeling. Ghosh et al. (Ghosh_Pargaonkar_Eisty_2024) evaluated fine-tuned LLMs for filtering non-informative user reviews on the Google Play Store and found them highly effective, achieving accuracies of 92% for both BERT and DistilBERT, and 91% for Gemma. Al Wahshat et al. (Al_Wahshat_Abu_ulbeh_Yusoff_Zakaria_Amir_Fazamin_Wan_Hamzah_P_2023) highlighted the potential of using GPT-4 in filtering manipulated reviews due to its powerful language processing capabilities.
Building on this line of work, our study integrates automated filtering as the first step of a unified, LLM-assisted framework that combines filtering with topic extraction and assignment. While prior work primarily evaluated filtering accuracy in isolation, we demonstrate that incorporating this process within a broader analysis framework improves data quality and supports large-scale review interpretation. Our filtering component achieves comparable accuracy (90%) while operating at a much larger scale across diverse Gen-AI app categories.
@@
Pre-LLM approaches to topic extraction. Before the advancement of LLMs, researchers used various methods to analyze user reviews. Genc Nayebi and Abran (Genc-Nayebi_Abran_2017) conducted a literature review in 2016 that explored approaches for mining opinions from user reviews. They identified three studies that relied on manual analysis, while others used automated techniques, such as Classification and Regression Trees (CART), word frequency statistics, correlation analysis, LDA, and Hexbin plots. Researchers and developers still use non-LLM approaches, even in recent work. For example, Arambepola et al. (10499727) used the Valence Aware Dictionary and sEntiment Reasoner (VADER) for sentiment analysis to categorize reviews into positive, neutral, and negative categories. They then used an LDA model for topic modeling of user reviews of educational apps. Their positive topics achieved a coherence score of 0.56, and their negative topics achieved a score of 0.49. Assi et al. (Assi_Hassan_Tian_Zou_2021) proposed the Global-Local sensitive Feature Extractor (GLFE) to identify high-level features in user reviews, achieving 79–82% precision and 74–77% recall on annotated review data.
These early techniques established the foundation for automated review analysis but relied heavily on statistical or lexicon-based models. With the advent of LLMs, topic extraction has shifted toward context-aware and semantically richer methods, as discussed next.
@@
LLM-based topic extraction and assignment methods. LLMs have demonstrated promising performance in tasks related to topic extraction and assignment. Arambepola et al. (Arambepola_Lalendra_Wimalasena_Munasinghe_2025) conducted a recent systematic review of mobile app review analysis techniques and found that integrating LLMs into app user review analysis provides a more holistic and context-rich approach to evaluating reviews and understanding users. Pham et al. (Pham_Hoyle_Sun_Resnik_Iyyer_2024) introduced TopicGPT, a framework to perform topic extraction by prompting Open AI’s GPT (openai2023gpt) to generate topics and assign them to documents. Their framework outperformed traditional topic models, such as LDA, SeededLDA, and BERTopic. Prakash et al. (Prakash_Wang_Hoang_Hee_Lee_2023) introduced PromptMTopics, a prompt-based ChatGPT model set to extract topics from textual meme captions that outperformed baseline models. While these studies evaluated LLM-based extraction on textual datasets such as documents or memes, our work applies similar techniques directly to large-scale user reviews. Assi et al. (Maram_LLM_Cure) introduced LLM-Cure, an approach that extracts and assigns features from user reviews to make suggestions for feature improvement. Their approach outperformed baseline models, achieving an F1-score of 85%. Among these, LLM-Cure most closely resembles our framework in combining both extraction and assignment tasks, though its focus was on generating feature suggestions rather than large-scale trend analysis.
@@
Single-task topic analysis. Instead of extracting and assigning topics to user reviews, some studies focused on a single task (i.e., either extracting topics from user reviews or assigning predefined topics to user reviews). For example, Wei et al. (Wei_Courbis_Lambolais_Xu_Bernard_Dray_2023) focused on review assignment using predefined labels rather than extracting topics. They assigned reviews using fixed categories such as feature requests, problem reports, and irrelevant feedback. While useful, this approach limits the discovery of emergent or evolving user concerns. Dos Santos et al. (Dos_Santos_Oliveira_De_Jesus_Aljedaani_Eler_2023) and Ren et al. (Ren_Nakagawa_Tsuchiya_2024) extracted topics from user reviews without performing the assignment task. While this approach is effective for uncovering topical insights, it does not account for the frequency or distribution of each topic, limiting its usefulness for trend or prevalence analysis.
Domain-specific applications of LLMs in user review analysis. Beyond general topic extraction, some studies applied LLMs to narrow domains of app analysis. Sorathiya and Ginde (Sorathiya_Ginde_2024) employed a hybrid approach combining Natural Language Inference (NLI) and LLMs to identify privacy concerns in mental health app reviews, uncovering ethically relevant feedback that keyword-based methods often miss. Rezaei Nasab et al. (RezaeiNasab_Dashti_Shahin_Zahedi_Khalajzadeh_Arora_Liang_2025) examined fairness concerns in AI-based mobile apps, using LLM and machine learning based classifiers to detect fairness mentions and cluster them into six categories.
While these studies reveal how LLMs can surface ethics-related insights within specific domains, our dataset and focus differ substantially. We analyze user reviews from a large-scale collection of Gen-AI apps spanning multiple functional categories, including but not limited to health & fitness, education, productivity, and entertainment. Unlike prior domain-specific datasets centered on single themes such as privacy or fairness, our corpus captures a broad range of emergent Gen-AI topics (i.e. AI performance, content quality, and content policy) that are unique to this new class of applications. This wider scope allows us to identify both category-specific differences and cross-domain similarities, showing how users across varied contexts often express comparable concerns and expectations about Gen-AI functionality.
Collectively, these studies demonstrate the growing potential of LLMs for review analysis yet reveal a gap in unified, scalable approaches that integrate filtering, topic extraction, and topic assignment across diverse app domains. To address this, we design the SARA framework, which operationalizes these steps into a cohesive process for large-scale analysis of Gen-AI user feedback.
@@ 3. Methodology
To study user feedback on Gen-AI apps, we propose and validate a structured four-phase framework we call SARA, which stands for Selection, Acquisition, Refinement, and Analysis. SARA encompasses 1) the selection of relevant Gen-AI apps, 2) the acquisition of user reviews, 3) a refinement data cleaning framework to ensure data quality for LLM analysis, and finally, 4) the LLM-based analysis of user reviews through topic extraction and assignment. SARA integrates and extends state-of-the-art methods in user review analysis (Maram_LLM_Cure; Chen_Lin_Hoi_Xiao_Zhang_2014), combining automated filtering, topic extraction, and topic assignment into a unified, empirically validated workflow that supports large-scale, LLM-driven analysis of user feedback. Although evaluated within the Gen-AI domain, SARA’s modular design enables its application to other mobile or AI-enabled systems for analyzing large-scale user feedback. We first evaluate the reliability of SARA in accurately labeling review topics (RQ1) and then apply it in subsequent analyses to explore user concerns and temporal trends (RQ2–RQ3). We present an overview of our research framework in Figure 2.
@@ 3.1. Selection of Gen-AI apps
    Keyword generation. We compile a list of keywords derived from a Gen-AI taxonomy by Gozalo-Brizuela and Garrido-Merchán (gozalobrizuela2023surveygenerativeaiapplications). The authors organized Gen-AI apps into 14 categories, each with a taxonomy of the relevant technology. We exclude four categories, i.e.,“biotech,” “brain,” “drug discovery,” and “other,” because an Android mobile app user does not typically use them. We create a list of ten categories relevant to mobile apps and repurpose the taxonomy as keywords, as shown in Table 1.
    Figure 2. Overview of our framework
@@
    Manual screening and validation. The initial search yields 303 apps. To ensure the inclusion of only genuine Gen-AI apps, the first and second authors independently review the apps by looking at the title, description, and images of the apps. We assess the inter-rater reliability, using Cohen’s Kappa (cohen1960kappa) ($\kappa=0.61$), which indicates substantial agreement. In case of disagreement, the two authors discuss and resolve the discrepancies for the ambiguous cases. This rigorous process results in a final list of 194 confirmed Gen-AI apps.
@@ 3.2. Acquisition of User Reviews
Following app selection, we collect user reviews using the Google Play Store Scraper (googleplayscraper). For each retrieved review, the crawler extracts a comprehensive set of metadata, including reviewId, userName, userImage, content (i.e., review text), score (i.e., user rating from 1 to 5 stars), thumbsUpCount, reviewCreatedVersion, at (i.e., timestamp of the review), replyContent, repliedAt, and appVersion. Our subsequent analysis primarily focuses on the content, score, and at fields. We organize the collected reviews based on the Google Play Store categories (app categories) of their respective apps. At the end of this step, we collect 2,778,643 reviews of 186 apps, as eight apps from our list have no textual reviews.
@@ 3.3. Multi-Stage User Review Refinement
To ensure the robustness and reliability of our LLM-based analysis, we implement a multi-stage data cleaning framework designed to enhance the quality of the user review data progressively. This multi-stage refinement process comprises four distinct stages. We refine data in stages because we are interested in experimenting with different levels of data cleaning in RQ1. We summarize the number of reviews retained after each cleaning stage in Table 2. Each stage builds upon the previous one as follows:
Table 2. Number of reviews by app category and stage (Full dataset considered).
@@
Stage 0: basic cleaning. Text preprocessing, as demonstrated in prior work (Ghosh_Pargaonkar_Eisty_2024; Prakash_Wang_Hoang_Hee_Lee_2023; Roumeliotis_Tselikas_Nasiopoulos_2024; Lee_Lee_Seo_2024), is a standard practice for eliminating noise and ensuring more effective analysis. It involves the removal of punctuation, special characters, extra white spaces, emojis, and irrelevant data such as usernames and URLs. In the initial stage, we exclude empty reviews and apps with no reviews, as we are interested in analyzing the textual content of the reviews. Then, we perform fundamental text preprocessing, including replacing commas with white spaces and removing special characters and emojis from the review content.
@@
We exclude five categories that contain fewer than 1,000 reviews, as small sample sizes tend to produce repetitive and low-quality topics. To validate this threshold, we experiment with categories of varying sizes and find that categories with fewer than 1,000 reviews yield topics that lack diversity and informativeness. For instance, the Personalization app category, which contains only 23 reviews, produces two nearly identical and non-informative topics labeled Positive Feedback and Praise. Additionally, some topics focus on non-informative content, such as Personal Comments or Signatures, which capture examples like “hello marlyn” and “Placau!!!!” that are not relevant to the app’s functionality or performance. Based on these observations, we adopt 1,000 reviews as the minimum threshold and exclude any app categories that fall below this cutoff.
@@
Stage 1: temporal filtering. This stage filters reviews written prior to 2022 to ensure Gen-AI relevancy. Some apps included in our initial screening are not originally designed as Gen-AI tools but have integrated Gen-AI functionalities at a later stage. For instance, in the case of Microsoft Bing Search [^4] https://play.google.com/store/apps/details?id=com.microsoft.bing, Bing Chat is a recently introduced feature that extends the capabilities of the search engine by incorporating Gen-AI models (Iorliam_Ingio_2024). Consequently, it is essential to exclude user reviews published before the integration of such features, as they reflect experiences with the app before the adoption of Gen-AI. While foundational work on AI models began as early as 2012, notable advancements in Gen-AI only started to gain momentum after 2018 (Sengar_Hasan_Kumar_Carroll_2024). The year 2018 also marked a rise in academic interest and publications related to Gen-AI (Dwivedi_Elluri_2024). However, the field witnessed a major breakthrough in 2022, particularly with the public release and widespread adoption of ChatGPT, which significantly accelerated the prominence and deployment of Gen-AI technologies across various platforms (Akhtar_2024). To maximize the relevance and accuracy of the user feedback analyzed in our study, we adopt 2022 as the cutoff year for including reviews.
@@
Stage 2: short review exclusion. Building upon the cleaned data from Stage 1, this stage employs a simple heuristic to further reduce noise by excluding reviews containing three words or less. Short reviews of three words or less do not contain context for meaningful topics (Assi_Hassan_Tian_Zou_2021; 10499727). This step aims to eliminate overly brief and potentially non-informative feedback.
@@
Stage 3: LLM-assisted informative review filtration. Non-informative reviews bring noise to user reviews analysis and influence the performance of topic extraction (Chen_Lin_Hoi_Xiao_Zhang_2014). To improve the quality of extracted topics, we exclude non-informative reviews. We define non-informative reviews as short, generic statements that express sentiment without referencing specific reasons or features. While Chen et al. (Chen_Lin_Hoi_Xiao_Zhang_2014) focused on informativeness from a developer’s perspective, we expand this definition to consider the informational needs of multiple stakeholders, including users, platform owners, developers and researchers. We carry out the LLM-assisted process in the following steps.
@@
    Prompt design. To filter non-informative reviews, we utilize Open AI’s (openai2023gpt) GPT-4o-mini model with zero temperature for reproducibility. We use gpt-4o-mini instead of the full gpt-4o model because it offers significantly faster response times while maintaining reasonable output quality for our task, which does not require complex reasoning (OpenAI_O4Mini_2024).
    Prior research comparing explicit, implicit, and creative prompting (Giray_2023; Ho_Mayberry_Nguyen_Dhulipala_Pallipuram_2024; Deshmukh_Raut_Bhavsar_Gurav_Patil_2025) shows that explicit, well-structured prompts produce more factual and task-aligned outputs, while underspecified ones invite unpredictable reasoning. Consistent with these findings and with approaches such as LLM-Cure (Maram_LLM_Cure), we design an explicit, structured prompt that classifies each review as either informative or non-informative. The prompt includes representative examples, a fixed binary label set, and a CSV-style output format to constrain the model’s responses. These measures ensure deterministic labeling and minimize hallucination by preventing free-form or extraneous text generation.
@@
    Iterative prompt refinement. Prior studies highlight that prompt structure and iterative refinement (“dynamic prompting”) substantially influence LLM performance  (Marvin_Hellen_Jjingo_Nakatumba-Nabende_2024; He_Rungta_Koleczek_Sekhon_Wang_Hasan_2024). In our setup, we adopt this principle by refining prompts through multiple pilot runs to improve label consistency and clarity. We use small statistically significant samples of 96 reviews (95% confidence interval, 10% margin of error) from three diverse app categories (Photography, Productivity, and Art & Design). Then, the first and second authors manually label the reviews in the samples as informative or non-informative. We calculate the inter-rater reliability using Kappa (cohen1960kappa) ($\kappa=0.91$), which indicates almost perfect agreement.
    Next, we design and test different versions of the filtering prompt against the manually labeled data, evaluating the model’s filtering accuracy and iteratively refining the prompt. Revisions include modifying the wording, structure, number of examples, and types of examples. We repeat this iterative process until we achieve a peak filtering accuracy of 91% on held-out validation samples. Figure 3 shows the final version of the filtering prompt, which includes ten examples (five informative and five non-informative) and yields optimal performance.